Can any one please help me!!!!
Evaluate the function for the value specified
f(x)=x^of 2 -7x+10; find (a)f(0 ),(b) f(5), and (c)f(-2)
Thank you so much for your help with this!
McGraw Hill 6th edition, chapter 7.5, exercise # 2

Answer by jenrobrody(19)

f(x)=x^of 2 -7x+10; find (a)f(0 ),(b) f(5), and (c)f(-2)
did you mean f(x)=x^2-7x+10?
if so:
a. f(0)=(0)^2 - 7(0) + 10
=0-0+10 = 10
b. f(5)= (5)^2- 7(5) + 10
= 25 - 35 + 10 = 0
c. f(-2)= (-2)^2-7(-2) + 10
= 4 + 14 + 10 = 28